# MFIT : Multi-FIdelity Thermal Modeling for 2.5D and 3D Multi-Chiplet Architectures

### v2

@@ III FEM for Thermal Analysis
FEM analysis begins by dividing the problem domain into small finite elements, converting the continuous governing partial differential equations (PDEs) into algebraic equations. Next, the system’s geometry is broken down into a lattice of small discrete cells called a “mesh” which approximates a larger, continuous block [28]. After applying the PDEs and boundary conditions to each element, the equations are assembled into a global algebraic system, maintaining continuity between adjacent elements. This global system, representing the discretized PDEs over the whole domain, is solved numerically for the field variables at each mesh node. In this work, only the equation governing solid conduction is solved [14]:
$\nabla\cdot\left[k\nabla T\right]+\dot{q}=\rho C_{v}\frac{\partial T}{\partial t}$ (1)
@@
III-A Stages of the FEM Simulation Pipeline
@@
Performing FEM simulations involves several key processing steps visualized in Figure 3. First, the geometry, a 3D representation of the 2.5D or 3D integrated package, is created using computer-aided design tools. This geometry should be as detailed as possible while allowing the setup and simulation to be completed within the given time constraints. Next, a volumetric mesh is generated by transforming the 3D model into one consisting of many individual cells on which the FEM software operates. This step can be iterated since creating appropriate mesh (aka meshing) is critical for computation time and accuracy of solutions. Once an acceptable mesh has been created, it is imported into the solver. The simulation is then set up, including boundary conditions, material parameters, power source terms, and other general model parameters. Our specific system setup is expanded upon in Section V. Finally, the FEM software simulates the model by solving the governing heat transfer equations.
III-B Impracticality of FEM Simulations
@@
While FEM simulations offer high accuracy, they are impractical for DSE or runtime thermal management due to their time-consuming setup and operation. The process of geometry creation, meshing, solver setup, and execution is intricate and often exceeds the simulation runtime itself. Because the simulation process requires multiple iterations for reliable results, this time overhead quickly becomes prohibitive. The setup of 2.5D or 3D integrated systems is especially complex due to the large number of discrete power sources. These systems also involve numerous small and large bodies, dramatically increasing the computational complexity and the solver runtime [28]. Simulation times range from hours to days, directly impacted by geometric detail, complexity, size, and setup parameters such as the time step. Consequently, analyzing 2.5D or 3D integrated systems with intricate geometries and operating conditions using FEM simulations becomes prohibitively time-consuming, highlighting the need for alternative approaches.
Figure 4: Flow diagram of MFIT, the proposed multi-fidelity thermal modeling framework.
Figure 5: Cross section of a 2.5D integrated system on Si-interposer showing abstracted blocks for the $\mu$-bumps, C4 bumps, and link structures.
Figure 6: Cross section of a 3D integrated system showing abstracted blocks for the $\mu$-bumps, C4 bumbs and link structures.
IV Multi-Fidelity Thermal Modeling
IV-A Overview of the Proposed Approach
@@
Our multi-fidelity thermal model set involves four individual models, visualized in Figure 4. The process of creating these models is identical to any packaging technology. MFIT considers 2.5D chiplet on silicon interposers and 3D direct bonded systems [9].
@@
We start by creating fine-grained FEM models of specific components within the package. For example, we model individual links within the interposer and $\mu$-bumps connecting a chiplet to the interposer. The cost of this level of detail is the model complexity and execution times that limit the simulation scope. Therefore, the fine-grained models are used as a reference to design abstracted FEM models, as explained in the following subsection. This abstraction enables system-level FEM simulations of systems with much higher chiplet counts than would otherwise be possible with negligible accuracy loss. All FEM simulations are performed using ANSYS Fluent [28]. The third class in the MFIT framework is the thermal RC models. Since these models are constructed using the geometry and material parameters of the system, a new RC model of a different system configuration can be created without re-running FEM simulations, allowing for rapid DSE. Finally, the continuous-time state-space equations that govern the thermal RC models are discretized with a given sampling period to create DSS models as detailed in Section IV-D.
IV-B Fine-Grained to Abstracted FEM Modeling
@@
MFIT focuses on two structures within a chiplet-based package for this work: the $\mu$-bumps connecting each chiplet to the interposer and the links that enable communication between chiplets. The rationale behind selecting these components is elaborated on in the following subsections. These two structures are present in both 2.5D and 3D chiplet-based packages, as shown in Figures 5 and 6, and the results of the abstraction experiments are applied to both the 2.5D and 3D full-system abstracted models.
@@
While we apply our abstract modeling approach to only two structures in this work, the same approach applies to other structures in the package, such as the substrate or C4 bumps. In addition to these abstractions, MFIT also models the heatsink as a heat transfer coefficient (HTC) instead of a physical model. This choice removes the need to model fluids in our simulations, as fluid flow is only used for convective heat transfer in the heatsink.
To capture the 3D chiplet-based systems, we consider direct die-to-die bonding between vertically stacked chiplets as an example. With this bonding method, no additional layers are modeled between stacked chiplets. Instead, chiplets are modeled directly contacting each other, stacked one on top of the other. Similarly, another bonding method can be utilized, such as TSVs or $\mu$-bumps between stacked chiplets. When modeling the geometry of TSVs or $\mu$-bumps between stacked chiplets in a system, an additional block of homogeneous material is created between each stacked chiplet. This block has the same thickness as the thickness of the bonding method. Then, to determine the material parameters of this new block, an identical process is followed to that described in the following section.
IV-B1 $\mu$-bump Abstracted Model
@@
The $\mu$-bumps are particularly important for thermal behavior since they are one of the two paths to dissipate heat away from a chiplet, as seen in Figure 5. Due to the density and the total number of $\mu$-bumps present, which number in the thousands, it is impractical to simulate an entire package with individually modeled $\mu$-bumps. Therefore, a small block of the $\mu$-bump layer, along with the associated chiplet and interposer, is simulated in isolation to determine thermal coefficients that can be applied to the final abstract models as illustrated in Figure 7.
Figure 7: Temperature contour of a $\mu$-bump layer subsection.
@@
First, the detailed block containing $\mu$-bumps and underfill material is simulated with static heat and convection boundaries, which are applied to create a measurable thermal gradient across the $\mu$-bump layer. Then, the thermal conductivity $k$ is calculated as:
$k=\frac{\dot{q}\cdot l}{A\cdot\Delta T}$ (2)
@@
where $\dot{q}$ is the heat flow rate, $l$ is the thickness of the material, $A$ is the cross-sectional area, and $\Delta T$ is the temperature difference across the material [40]. Thermal capacitance and specific heat are calculated via weighted body average [41]. These parameters are applied to a model containing a homogeneous block in place of the previously modeled $\mu$-bumps and underfill material. Finally, the same boundary conditions are used, as shown in Figure 7.
@@
We observe identical temperature drop across the $\mu$-bump layer of the abstracted model and less than a tenth of a degree difference in interface temperatures in this sub-block, as presented in Table I, while achieving approximately 1.5x speedup.
@@
TABLE I: Temperature results of the $\mu$-bump block abstraction experiments. Only a single result is shown for brevity.
@@
IV-B2 Link Abstracted Model
A link is a group of wires embedded in the interposer to interconnect chiplets. Depending on the thermal crosstalk over the links, the NoI architecture can play a significant role in a system’s thermal behavior. To determine how links are modeled in the complete system simulations, we tested three different configurations of a two-chiplet package. These configurations model the link (1) in full detail, (2) as an abstracted block, and (3) not model at all. We use two different power configurations: (a) the power dissipation is static over time, and (b) it varies dynamically over time. These power consumption profiles are applied to one chiplet while the temperature of the other chiplet is calculated through FEM simulation. The mean absolute error (MAE) of the receiving chiplet temperature compared to the detailed model case is recorded in Table II. Only a minimal accuracy loss is observed for both cases, while execution time savings are significant, as shown in Table III. Therefore, we choose not to model links in our full-system simulations.
TABLE II: Error comparison between abstracting and removing the links compared to detailed link modeling.
@@ IV-B2 Link Abstracted Model
Steady
@@
MAE (C)
Abstracted links 0.05 0.02
No links 0.34 0.13
TABLE III: Execution time for Detailed, Abstract, and No Links experiments with steady and transient power inputs.
@@
IV-B3 Heatsink Abstracted Model
@@
FEM simulations that involve a heatsink must model the convective heat transfer from the heatsink to the atmosphere using fluid models [30]. However, modeling fluid dramatically increases the setup and simulation time. Additionally, the geometry must be modified for every heatsink configuration, further increasing the time needed for design iteration. Due to their complexity, high-performance cooling methods, such as liquid cooling, are also difficult to model in FEM-based simulations.
@@
MFIT removes the need to model the heatsink by abstracting the cooling solution to a single HTC. We then apply this coefficient to the top of the lid where a heatsink would typically sit. Modeling a heatsink as an HTC is an active area of research that has been studied for many different cooling solutions [42, 43]. This approach allows for a great deal of flexibility in FEM modeling. Instead of completing the time-consuming pipeline in Figure 3, the HTC can be modified to test the behavior of different cooling solutions.
@@
MFIT, assumes an active air-cooled heatsink. The value of the HTC of an air-cooled heatsink is determined by:
$h_{eq}=\frac{h_{avg}\cdot A_{t}\cdot\left(1-\frac{N\cdot A_{f}\cdot(1-\eta_{f})}{A_{t}}\right)}{LW}$ (3)
@@
where $A_{t}$ is the total area of the heatsink, $A_{f}$ is the fin area, $N$ is the number of fins, $\eta_{f}$ is the fin efficiency, and $L$ and $W$ are the length and width of the base plate. While the average convective HTC ($h_{avg}$) can be calculated using the Nusselt number [41]. We select values consistent with a basic copper heatsink with forced airflow provided by a typical commercial computer fan. The parameters can be easily tuned to reflect different cooling solutions. Applying this method accelerates the simulation process substantially.
@@
While this method is effective for systems where heat is dissipated primarily through the lid, a different approach may be required for other cooling methods, such as inter-tier liquid cooling, where microchannels contact the chip directly [39, 38]. In such an approach, heat is dissipated directly from the chip without moving to a heat spreader like the lid. Hence, additional heat transfer coefficients or abstraction techniques may be needed to capture the cooling behavior.
IV-C FEM to Thermal RC models
TABLE IV: Summary of the key parameters used in this work.

### v4

@@ 3. Finite Element Method for Thermal Analysis
FEM analysis begins by dividing the problem domain into small finite elements, converting the continuous governing partial differential equations (PDEs) into algebraic equations. Next, the system’s geometry is broken down into a lattice of small discrete cells called a “mesh” which approximates a larger, continuous block (Matsson, 2023). After applying the PDEs and boundary conditions to each element, the equations are assembled into a global algebraic system, maintaining continuity between adjacent elements. This global system, representing the discretized PDEs over the whole domain, is solved numerically for the field variables at each mesh node. In this work, only the equation governing solid conduction is solved (Sultan et al., 2019):
(1) $\nabla\cdot\left[k\nabla T\right]+\dot{q}=\rho C_{v}\frac{\partial T}{\partial t}$
@@
3.1. Stages of the FEM Simulation Pipeline
@@
Performing FEM simulations involves several key processing steps visualized in Fig. 3. First, the geometry, a 3D representation of the 2.5D or 3D integrated package, is created using computer-aided design tools. While increasing the detail in the geometry allows for more accuracy in modeling the system, it also increases the simulation time. This geometry should be as detailed as possible while allowing the setup and simulation to be completed within the given time constraints. Next, a volumetric mesh is generated by transforming the 3D model into one consisting of many individual cells on which the FEM software operates. A mesh sensitivity study is performed to determine if the mesh quality is sufficient. In this process, simulations are run with a progressively finer mesh while output parameters (such as temperature) are monitored. The mesh quality is considered to be sufficient when increasing the mesh granularity no longer impacts the temperature output.
Once an acceptable mesh has been created, it is imported into the solver. The simulation is then set up, including boundary conditions, material parameters, power source terms, and other general model parameters. Our specific system setup is expanded upon in Section 5. Finally, the FEM software simulates the model by solving the governing equations.
Figure 3. An illustration of the FEM pipeline.
3.2. Impracticality of FEM Simulations
@@
While FEM simulations offer high accuracy, they are impractical for DSE or runtime thermal management due to their time-consuming setup and operation. The process of geometry creation, meshing, solver setup, and execution is intricate and often exceeds the simulation runtime itself. Because the simulation process requires multiple iterations for reliable results, this time overhead quickly becomes prohibitive. The setup of 2.5D or 3D integrated systems is especially complex due to the large number of discrete power sources. These systems also involve numerous small and large bodies, dramatically increasing the computational complexity and the solver runtime (Matsson, 2023). Simulation times range from hours to days, directly impacted by geometric detail, complexity, size, and setup parameters such as the time step. Consequently, analyzing 2.5D or 3D integrated systems with intricate geometries and operating conditions using FEM simulations becomes prohibitively time-consuming, highlighting the need for alternative approaches.
4. Multi-Fidelity Thermal Modeling
4.1. Overview of the Proposed Approach
@@
Our multi-fidelity thermal model set involves four individual models, visualized in Fig. 4. The process of creating these models is identical for any packaging technology. In this work, we consider 2.5D chiplet on silicon interposers and 3D systems integrated with $\mu$-bumps (Agarwal et al., 2022; Lau, 2023). MFIT supports other integration methods and packaging technologies as well, as described below.
Figure 4. The proposed workflow process to produce the multi-fidelity set of thermal models, starting with the most accurate yet slow models and deriving faster models.
@@
We start by creating fine-grained FEM models of specific components within the package. For example, we model individual links within the interposer and $\mu$-bumps which are used to connect chiplets to the interposer and between 3D stacked chiplets. The cost of this level of detail is the model complexity and execution time that limits the simulation scope. Therefore, these fine-grained models are used as a reference to design abstracted FEM models, as explained in the following subsection. This abstraction enables system-level FEM simulations of systems with significantly higher chiplet counts than would otherwise be possible, with negligible accuracy loss. All FEM simulations are performed using ANSYS Fluent (Matsson, 2023). The third class of model in the MFIT framework is the thermal RC model. Since these models are constructed using the geometry and material parameters of the system, a new RC model of a different system configuration can be created without re-running FEM simulations, allowing for rapid DSE. Finally, the continuous-time state-space equations that govern the thermal RC models are discretized with a given sampling period to create DSS models as detailed in Section 4.4.
4.2. Fine-Grained to Abstracted FEM Modeling
@@
MFIT focuses on two structures within a chiplet-based package for this work: the $\mu$-bumps connecting each chiplet to the interposer and 3D stacked chiplets to each other and the links that enable communication between chiplets. The rationale behind selecting these components is elaborated on in the following subsections. These two structures are present in both 2.5D and 3D chiplet-based packages, as shown in Figs. 5 and 6, and the results of the abstraction experiments are applied to both the 2.5D and 3D full-system abstracted models.
@@
While we apply our abstract modeling approach to only two structures in this work, the same approach applies to other structures in the package, such as the substrate or C4 bumps. In addition to these abstractions, MFIT also models the heatsink as a heat transfer coefficient (HTC) instead of a physical model. This choice removes the need to model fluids in our simulations, as fluid flow in our simulations is used only for convective heat transfer in the heatsink.
Next, we present our abstraction techniques and their validation instead of leaving them to the experimental evaluation to improve the readability and to justify larger-scale experiments performed using the abstract models.
Figure 5. Cross section of a 2.5D integrated system on Si-interposer, showing abstracted blocks for the $\mu$-bumps, C4 bumps, and link structures.
Figure 6. Cross section of a 3D integrated system, showing abstracted blocks for the $\mu$-bumps, C4 bumps and link structures.
4.2.1. $\mu$-bump Abstracted Model
@@
The $\mu$-bumps are particularly important for thermal behavior due to their placement in the 2.5D and 3D integrated systems. In the 2.5D system, the $\mu$-bumps are one of two paths to dissipate heat from a chiplet, as illustrated in Fig. 5. Similarly, in the 3D system, all heat flow for the lower two chiplets in each stack must travel through one of two $\mu$-bump layers, as shown in Fig. 6. Due to the density and the total number of $\mu$-bumps present, which number in the thousands, it is impractical to simulate an entire package with individually modeled $\mu$-bumps in FEM. Therefore, a small block of the $\mu$-bump layer, along with the associated upper and lower layer, is simulated in isolation to determine thermal properties that can be applied to the final abstracted models as illustrated in Fig. 7.
Figure 7. Temperature contour of a $\mu$-bump layer subsection.
@@
First, the detailed block containing $\mu$-bumps and underfill material is simulated with static heat flux and convection boundaries, which are applied to create a measurable thermal gradient across the $\mu$-bump layer. Then, the thermal conductivity $k$ is calculated as:
(2) $k=\frac{\dot{q}\cdot l}{A\cdot\Delta T}$
@@
where $\dot{q}$ is the heat flow rate, $l$ is the thickness of the material, $A$ is the cross-sectional area, and $\Delta T$ is the temperature difference across the material (Nellis and Klein, 2008). Thermal capacitance and specific heat are calculated via weighted body average (L. Bergman et al., 2011). These parameters are applied to a model containing a homogeneous block in place of the previously modeled $\mu$-bumps and underfill material. Finally, the same boundary conditions are used, as shown in Fig. 7.
@@
We observe identical temperature drop across the $\mu$-bump layer of the abstracted model and less than a tenth of a degree difference in interface temperatures in this sub-block, as presented in Table 2, while achieving approximately 1.5x speedup.
@@
Table 2. Temperature results of the $\mu$-bump block abstraction experiments. Only a single result is shown for brevity.
@@
4.2.2. Link Abstracted Model
A link refers to a group of wires embedded in the interposer that interconnect chiplets. Depending on the thermal crosstalk through these links, the Network-on-Interposer (NoI) architecture can influence the overall thermal behavior of a system. To understand the significance of accurately modeling links in full-system thermal simulations, we evaluated three configurations for a two-chiplet package:
1. Full-detail link model: This configuration explicitly models each horizontal and vertical wire segment in the interposer layer, including their placement and density between the chiplets. Specifically, we implemented two $\times$64 Advanced Package Lane interfaces, following the bump map layout described in the UCIe specification (Sharma et al., 2022b). This detailed model captures the realistic routing and thermal characteristics of Redistribution Layer (RDL)-based interposer connections.
2. Abstracted block model: Here, the link is represented as a single thermal block with averaged properties, simplifying its spatial granularity.
3. No-link model: In this case, the link is completely omitted from the thermal model.
We use two different power configurations: (a) the power dissipation is static over time, and (b) it varies dynamically over time. These power consumption profiles are applied to one chiplet while the temperature of the other chiplet is calculated through FEM simulation. The mean absolute error (MAE) and average percentage error of the receiving chiplet temperature compared to the detailed model case are recorded in Table 3. Only a minimal accuracy loss is observed for both cases, while execution time savings are significant, as shown in Table 4. Therefore, we choose not to model links in our full-system simulations.
Table 3. MAE and Average Percentage Error comparison between abstracting and removing the links compared to detailed link modeling while operating between 60C and 100C.
@@ 4.2.2. Link Abstracted Model
Steady-State
MAE (C)
Steady-State
Avg. % Error
Transient
@@
Avg. % Error
Abstracted links 0.05C 0.083% 0.02C 0.045%
No links 0.34C 0.429% 0.13C 0.259%
Table 4. Execution time for Detailed, Abstract, and No Links experiments with steady and transient power inputs.
@@
4.2.3. TSV Abstracted Model
Silicon interposers with through silicon vias (TSVs) are one of the prominent methods for integrating multiple chiplets into a single package. In silicon interposers and 3D stacked chips, TSVs are crucial structures to conduct signals and power through the chip to layers above and below. We develop and validate abstracted TSV models similar to the $\mu$-bump modeling presented in Section 4.2.1.
The first step is modeling a detailed block of the TSV layer, including the TSVs themselves and the surrounding silicon material. Additional blocks are placed above and below this layer to represent adjacent layers present in the full system. To model the thermal conductivity of the TSV later, heat flux and convection are applied to the top and bottom blocks, respectively. This standard process (Nellis and Klein, 2008) a temperature difference $\Delta$T across the TSV layer, as visualized in Fig. 7 for the u-bump layer. Then, Equation 2 is used to find the effective thermal conductivity $k$.
The second step is to create the abstracted thermal model, where the TSV layer is converted from a detailed representation to a single homogeneous block. The effective thermal conductivity found using the detailed model above is applied to this homogeneous block. Then, the same heat flux and convection which were applied to the detailed model are applied to this abstracted geometry. Finally, the temperature difference is measured again at the top and bottom of the TSV layer and compared to the temperature difference found using the detailed model.
[…]
Table 5. Temperature results of the TSV block abstraction experiments. Results are shown
TSV type Model
Top surface
temperature (C)
Bottom surface
Temperature (C)
Temperature
Drop (C)
[…]
Abstracted 84.77 61.18 23.59
[…]
Abstracted 85.87 61.18 24.69
4.2.4. Heatsink Abstracted Model
@@
FEM simulations that involve a heatsink must model the convective heat transfer from the system to the atmosphere using fluid models (Zhou et al., 2022). However, modeling the fluid dynamics dramatically increases the setup and runtime of FEM simulations. Additionally, the geometry must be modified for every different heatsink configuration, further increasing the time needed for design iteration. Due to their complexity, high-performance cooling methods, such as liquid cooling, are also difficult to model in FEM-based simulations.
@@
In MFIT, we remove the need to model the heatsink by abstracting the cooling solution to a single HTC. This coefficient is applied to the top of the lid where a heatsink is typically attached. Modeling a heatsink as a HTC is a common practice for many different cooling solutions (Khan et al., 2008; Narasimhan and Majdalani, 2002). This approach allows for a great deal of flexibility in FEM modeling. Instead of completing the time-consuming pipeline in Fig. 3, the HTC can be easily modified to test the behavior of different cooling solutions.
@@
The experiments demonstrated in this work assume an active air-cooled heatsink. The value of the HTC of an air-cooled heatsink is determined by:
(3) $h_{eq}=\frac{h_{avg}\cdot A_{t}\cdot\left(1-\frac{N\cdot A_{f}\cdot(1-\eta_{f})}{A_{t}}\right)}{LW}$
@@
where $A_{t}$ is the total area of the heatsink, $A_{f}$ is the fin area, $N$ is the number of fins, $\eta_{f}$ is the fin efficiency, and $L$ and $W$ are the length and width of the base plate. The average convective HTC ($h_{avg}$) can be calculated using the Nusselt number (L. Bergman et al., 2011). We select values consistent with a basic copper heatsink with forced airflow provided by a typical commercial computer fan.
While an active air-cooled heatsink is assumed in this work, similar equations to Equation 3 for liquid cooling exist (L. Bergman et al., 2011). If the system designer chose to use liquid cooling for the simulation, a new HTC would be calculated and applied in place of the air cooling HTC described previously.
@@
While the above method of abstracting a heatsink is effective for systems where heat is dissipated primarily through the external boundaries of the package (top of the lid, bottom of the substrate), a different modeling approach may be used for other cooling methods, such as inter-tier liquid cooling, where microchannels contact the chip directly (Hwang et al., 2018; Kwon et al., 2020). In such an approach, heat is dissipated directly from the chip without moving to a heat spreader like the lid. Hence, additional heat transfer coefficients or abstraction techniques may be needed to capture the cooling behavior.
4.3. FEM to Thermal RC models
This section describes the process of constructing a thermal RC model from the geometry of a given package. MFIT applies this technique to both 2.5D and 3D systems demonstrating the flexibility of the proposed methodology to different packaging technologies. However, this process can easily be applied to any package. Table 6 provides the notation of the equations used in this and the following subsection.
Table 6. Summary of the key parameters and notations used in this work.